Week of 19 March 2018

Problem 12.42

5 of 7

A 4-ft-diameter unfinished-concrete drain pipe has been designed to transport 90 ft3/s of water at a depth of 3 ft .

Part A

Determine the required downward slope of the pipe.

Express your answer using three significant figures.

S0=

.

SubmitPrevious AnswersRequest Answer

Incorrect; Try Again; 5 attempts remaining

Provide Feedback

Next

Incorrect. Incorrect; Try Again; 5 attempts remaining. No additional feedback.
Answer #1

solution :- im T 3m &=am for infinshed concrete n=0.017 e = cost ( 1 ) = 7 & = 27-2XT/2 = 45 A= R2 (B - SmB) = 22 (UT - SMAR)